Solve each of the following problems. Be sure to show the equations used for the solution.
Number problems. Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?

Answer by ankor@dixie-net.com(12701)


Write an equation for each sentence:
:
"Jill has $3.50 in nickels and dimes."
.05n + .10d = 3.50
:
"If she has 50 coins,"
n + d = 50
n = 50 - d; use for substitution in the 1st equation
:
how many of each type of coin does she have?
.05(50-d) + .10d = 3.50
2.5 - .05d + .10d = 3.50
-.05d + .10d = 3.50 - 2.5
+.05d = 1.00
d = 1/.05
d = 20 dimes
:
you can find the number of nickels using the equation: n + d = 50
:
Check the solutions by substitution in: .05n + .10d = 3.50